# Information Theory for Complex Systems Scientists: What, Why, & How?

### v3

@@ II.1.4 Conditional Entropy
The conditional entropy quantifies how uncertain we are about the state of variable given that we have knowledge of another variable. As with many information theoretic measures, it is best understood in a Bayesian context: our uncertainty about a variable will change depending on what information we have about the context around it. The entropy of a variable describes uncertainty about the state of a variable and is not a value intrinsic to elements of the real world in the way something like mass is. Our level of certainty changes depending on what information we have.
@@ II.3 Mutual Information
Like the Shannon entropy, the mutual information between two variables $X_{1}$ and $X_{2}$ is one of the foundational measures of information theory. Mutual information provides a measure of the statistical dependency between two random variables: how much does knowing the state of $X_{1}$ reduce our uncertainty about the state of $X_{2}$ (and vice versa). Recall from Sec. II.1.2 that “information” is equivalent to “reduction in uncertainty”, so mutual information provides the first real measure of information we have encountered so far. For two random variables, it has a number of equivalent forms:
@@ II.3.2 Conditional mutual information
There are two contexts in which this can occur: when information is redundantly shared over multiple variables and when information is synergistically present in the joint states of multiple variables. In both cases, examining only the pairwise relationships between elements provides an inaccurate picture of the system under study. Generally, redundant information will result in over-estimates of the extent to which elements are interacting, while synergistic information will result in under-estimates of th same. As we will discuss in Sections V.1 and V.2, this can significantly complicate the problem of constructing network models of complex systems from data.
@@ II.4 Multivariate Generalization of Mutual Information
Figure 2: This Venn diagram generalizes the relationships introduced in Figure 1. However, care should be taken when considering entropy diagrams for more than two variables: the innermost intersection ($I(X_{1};X_{2};X_{3})$, corresponding to the co-information or interaction information) is not strictly positive [matsuda_physical_2000]. The sign refers to whether the trivariate relationship is redundancy or synergy dominated [61], discussed below.
A defining feature of complex systems, however, is that they are comprised of a large number of interacting elements, often numbering in the hundreds or thousands. This has lead to a natural interest in information theoretic tools that can assess the structure of many-element systems and how they can interact. Arguably the most straight-forward attempts have been multivariate generalizations of mutual information, although this turns out to not be as straightforward as one might think.
@@ II.4.1 Total Correlation
$TC(\textbf{X})=D_{KL}(P(X_{1},X_{2},...X_{N})||P(X_{1})\times P(X_{2})...P(X_{N}))$ (39)
@@ II.4.3 Co-Information
Where $\xi$ indicates all possible subsets of X. Unfortunately, unlike the total and dual total correlations, which easily accommodate large numbers of variables without changing the interpretation, the co-information can be hard to interpret for more than 3 elements. A significant issue is that the co-information can be negative, which has historically limited it’s applicability outside of theoretical contexts. [61] showed that, for three variables, the sign of co-information can be understood as indicating whether the statistical dependencies within the dataset are dominated by redundancy or synergy, although this pattern does not necessarily generalize (compare with the O-information, introduced in Sec. VI).
@@ III Information Dynamics
To begin with, what does it mean for a system to “compute?” This inevitably raises the complex, philosophical question “what is computation”, which we will sidestep entirely by operationalizing “computation” in a complex system as the process by which the system decides its immediate next state. This can happen at every scale: individual elements can be in multiple states and decide their immediate next state based on the interactions between their own internal dynamics, as well as all the other elements that our single element is causally connected to. The joint-state of the whole system makes up the macro-state of the whole which “emerges” from the interactions of all it’s constituent elements [64].
@@
    Information storage [65], which characterizes how the past of a single element constrains its immediate future,
@@ III.1 Conditional Entropy Rate
where $k$ indicates the number of discrete past states to take into account. Following the usual formalism, $X_{t}$ refers to the state of $X$ at time $t$, while $X_{t-1}^{-k}$ indicates a vector of states, stretching from time $t-1$ all the way back to time $t-1-k$. This tells us how surprising any particular realization $x$ is, given that it is preceded by the particular sequence $x_{-k:t-1}$. When $h_{\mu}(x)$ is high, then X has done something very unexpected and inconsistent with its past behavior, while when $h_{\mu}(x)$ is low, whatever state $X$ has just adopted would be unsurprising to anyone who had been watching it for the last $k$ timesteps. The local conditional entropy rate has been used in thermodynamic interpretations of the transfer entropy measure [74]. The local conditional entropy rate can also been used to “whiten” an autocorrelated signal [75, 76]. Using this technique, a time series is transformed by replacing the value at every frame with the local conditional entropy rate of that frame, resulting in a time series with no autocorrelation.
@@ III.2 Information Storage
The conditional entropy quantifies the ”left over” uncertainty intrinsic to $X$ after the past has been accounted for. Its dual measure is the active information storage [65, 72], which quantifies how much information the past of a variable discloses about its immediate next state. In this context, “information storage” can be thought of as a measure of how dependent an element of a system is on it’s own past. If every next state is chosen at random, then $H_{\mu}(X)=H(X)$ and $A(X)=0$. Formally, active information storage is given by:
@@
Intuitively, this shows us that some of our uncertainty about the next state of $X$ could be resolved by looking at $X$’s past, and a residual part that cannot be resolved. This is a key insight when attempting to model complex systems. For processes with a naturally high entropy rate, simply observing the past behavior will not help predict the future (and consequently, modeling is hard). Active information storage has been applied in a wide range of contexts, from understanding distributed processing in neural networks [72], predictive coding models [77], eye and gaze tracking [78], to understanding the dynamics of football matches [79].
@@
This allows a “time-resolved” picture of how a single variable remembers and computes with it’s own past. When $a(x)>0$, the system is behaving “as we would expect”, in that past states are reducing our uncertainty about future states in accordance with the overall statistics of the system. In contrast, when $a(x)<0$, we are being “misinformed” and the system appears to be transiently “breaking” from it’s own past to do something unexpected.
@@ III.3 Information Transfer
Transfer entropy is probably the most widely used information dynamic in the study of complex systems, having been used in studies of biological networks of neurons, [85, 86, antonello_self-organization_2022, 87, 88, 89], cryptocurrencies [90, 91, 92, 93], precious metal economics [94], the interaction between social media sentiment and commodity prices [93], the directed relationship between atmospheric carbon dioxide levels and global warming [95], climate models [96, 30, 97] and more (for further citations, see [84]). Transfer entropy’s nonparametric nature makes it particularly suited to analysis of complex systems where non-linearities are known to play a significant role in the dynamics.
@@ III.3.1 Conditional Transfer Entropy
The effect of conditioning on $Z$’s past gives us the amount of information $X$’s past gives us about $Y$’s future above and beyond that which is provided by $Y$’s own past and $Z$’s past. As with conditional mutual information, conditional transfer entropy is essential when assessing information dynamics in complex systems, which by definition have large numbers of interacting elements. For two variables $X$ and $Y$ belonging to a set of elements V, the complete transfer entropy between $X$ and $Y$ ($TE(X\to Y|\textbf{V})$) is given by conditioning $TE(X\to Y)$ on the past of every other element in V simultaneously. The conditional transfer entropy plays a key role in the discovery of effective connectivity networks from time-series data (see Sec. V.2 and [106, 107]). Conditional transfer entropy can be localized in the same way as the standard bivariate transfer entropy.
@@ IV.1 PID Intuition
Where $Red(X_{1},X_{2};Y)$ corresponds to the information about $Y$ that is redundantly present in both $X_{1}$ and $X_{2}$ (i.e. could be learned by observing either $X_{1}$ or $X_{2}$.). $Unq(X_{1};Y/X_{2})$ is information about $Y$ that is uniquely present in $X_{1}$ (i.e. information that could only be learned by observing $X_{1}$), and likequise for $Unq(X_{2};Y/X_{1})$. Finally, $Syn(X_{1},X_{2};Y)$ is that information about $Y$ that is synergistically present in the joint states of $X_{1}$ and $X_{2}$ (and could not be learned by examining either $X_{1}$ or $X_{2}$ individually). A moments reflection should convince the reader that this enumerates all possible combinations.
@@
In general, however, for $N>2$, it is not the case that the relationship between mutual information terms and partial information “atoms” is so tightly constrained. For $N=3$, there are seven known mutual information terms but eighteen unique partial information atoms. To solve the general care involves developing some mathematical machinery beyond that which is provided by classical Shannon information.
@@ IV.2 Redundancy Lattices & Functions
Suppose we have a system of $N$ predictor variables, all informing about the state of a single target $Y$. We would like decompose the joint mutual information $I(X_{1},\ldots,K_{N};Y)$ down into all relevant combinations of sources. Doing this requites two things: the first is an operational definition of what it means for information to be “redundant” to two or more sources (which we will denote as $I_{\cap}$). The second is an understanding of the structure of multivariate information (which is given by the redundancy lattice described below) and Formally, given some (as yet undefined) set of all meaningful sources of information ($\mathfrak{A}$), we want a decomposition such that:
@@
Where $I_{\partial}$ quantifies the information about $Y$ that could only be learned by observing $\boldsymbol{\alpha}$ and no simpler combination of sources.
@@ IV.2.1 The Redundancy Lattice
When decomposing $I(X_{1},\ldots,X_{N};Y)$, the first step is to enumerate the set of all possible sources of information. For example, information can be disclosed by any $X_{i}$, but it can also be disclosed by any pair {$X_{i},X_{j}\}$, every triple, and so on. The set S of all possible (potentially multivriate) sources is given by the power set of the set of $\{X_{i}\}$: $\textbf{S}=\mathcal{P}_{0}(\{X_{1},\ldots,X_{N}\})$, where the 0 subscript indicates that the empty set is being excluded from the power set.
@@
The foure elements of $\mathfrak{A}$ correspond to our intuitive ideas of redundant, unique, and synergistic information described above. The partial ordering shows that $\{X_{1}\}\{X_{2}\}\prec\{X_{1}\}$ and $\{X_{2}\}$, while both $\{X_{1}\}$ and $\{X_{2}\}\prec\{X_{1},X_{2}\}$. This defines the redundancy lattice shown in Fig. 4.
@@ IV.3 Information Modification, Revisited
This shows that the transfer entropy consists of two, non-overlapping informational components. The first is referred to as the state-independent information transfer: $Unq(X_{-k:t-1};Y_{t})$ is the information that passes from $X$’s past of $Y$’s future and is independent of the state of $Y$’s own past. In contrast, the state-dependent information transfer, $Syn(X_{-k:t-1},Y_{-j:t-1};Y_{t})$, which is the novel information generated at time $t$ in $Y$ by the interaction of $X$’s past and $Y$’s past.
This complicates transfer entropy’s status as a measure of information transfer, as it appears to conflate the true transfer of information with synergistic information modification [134]. As it stands, the significance of this has largely remained under-explored. Daube, Gross, and Ince explored the contribution of synergy to sensory EEG data analysis, finding that the apparent transfer entropy was largely reflecting the synergistic component (with the unique “true flow” component being negative) [75]. This is one area in particular with a lot of potential to support future work.
@@ IV.4.3 Integrated Information Decomposition
which can be computed for continuous variables under Gaussian assumptions, but does not readily localize to individual realizations. Unfortunately, while the $I^{MMI}_{\Phi\cap}$ function has practical appeal, it has counter-intuitive behaviors that have led some to question its appropriateness [148]. In particular, it does not appears to conflate first-order autocorrelation and higher-order temporal synergy in potentially fatal ways. Other double redundancy functions have been proposed, including one generalizing the $i^{sx}_{\cap}$ function [145], and one based on the $i^{ccs}_{\cap}$ function [70], and one based on the dependency lattice approach of James et al., [119, 69].
@@ V Network Inference
One of the foundational methods in complex systems science is the construction of network models from data [152, 153]. It is hard to overstate how fundamental networks are to the modern study of complex systems, appearing in everything from brain science [154] to [155, 156], from ecology [157, 158] to economics [159]. In applications, networks generally break down into two broad classes: there are physical networks (where the nodes and edge are unambiguous, physical objects in the real world), and statistical networks (where the links between nodes reflects some statistical dependency between them). Physical networks are generally reasonably easy to infer the structure of. For example, the airline network is conceptually simple: nodes may be airports (or cities), and a directed edge exists between two nodes if there is a regular flight from one to another [160]. Another example might be a structural connectivity network map of the brain: using diffusion tensor imaging (DTI) it is possible to get an estimate of the white matter tracts in the brain that physically connect disparate brain regions [161].
In contrast, statistical networks are more abstract: rather than referring directly to some physical structure in the real world (like a white-matter tract, or airline route), the edges in a statistical network refer to some degree of statistical dependency between them. In an information-theoretic context, these dependencies almost always take the form of some kind of mutual information: nodes are usually said to be connected if knowing something about the state of one node reduces our uncertainty about the state of another node. There are other “kinds” of statistical networks not based on mutual information (for example, dynamic causal modeling builds statistical networks using Bayesian model comparison [162, 163]), however for our purposes, we will focus on information-theoretic approaches. Within the category of statistical networks, we will describe two general classes: the “functional” connectivity network, which has no notion of temporal order and produces undirected networks, and the “effective” connectivity network, which does take time into account (if time is a relevant feature) and produces directed networks. Finally, we will discuss how the problem of synergies in data can complicate the very idea of a bivariate network model of a complex system and briefly touch on higher-order models such as hypergraphs and simplicial complexes (although a rigorous mathematical treatment of these topics is beyond the scope of this tutorial).

### v4

@@ II.1.4 Conditional Entropy
The conditional entropy quantifies how uncertain we are about the state of variable given that we have knowledge of another variable. As with many information theoretic measures, it is best understood in a Bayesian context: our uncertainty about a variable will change depending on what information we have about the context around it. The entropy of a variable describes your uncertainty about the state of that variable, and is not a value intrinsic to the real world in the way something like mass is. Our level of certainty changes depending on what information we have.
@@ II.3 Mutual Information
Like the Shannon entropy, the mutual information between two variables $X_{1}$ and $X_{2}$ is one of the foundational measures of information theory. Mutual information provides a measure of the statistical dependency between two random variables: how much does knowing the state of $X_{1}$ reduce our uncertainty about the state of $X_{2}$ (and vice versa)? Recall from Sec. II.1.2 that “information” is equivalent to “reduction in uncertainty”, so mutual information provides the first real measure of information we have encountered so far. For two random variables, it has a number of equivalent forms:
@@ II.3.2 Conditional mutual information
There are two contexts in which this kind of three-way interaction can occur: when information is redundantly shared over multiple variables and when information is synergistically present in the joint states of multiple variables. In both cases, examining only the pairwise relationships between elements provides an inaccurate picture of the system under study. Generally, redundant information will result in over-estimates of the extent to which elements are interacting, while synergistic information will result in under-estimates of th same. As we will discuss in Sections V.1 and V.2, this can significantly complicate the problem of constructing network models of complex systems from data.
@@ II.4 Multivariate Generalization of Mutual Information
Figure 2: This Venn diagram generalizes the relationships introduced in Figure 1. However, care should be taken when considering entropy diagrams for more than two variables: the innermost intersection ($I(X_{1};X_{2};X_{3})$, corresponding to the co-information or interaction information) is not strictly positive [52]. The sign refers to whether the trivariate relationship is redundancy or synergy dominated [53], discussed below.
A defining feature of complex systems, however, is that they are comprised of a large number of interacting elements, often numbering in the hundreds or thousands. This has lead to a natural interest in information theoretic tools that can assess the structure of many-element systems and how they can interact. Arguably the most straightforward attempts have been multivariate generalizations of mutual information, although this turns out to not be as straightforward as one might think.
@@ II.4.1 Total Correlation
$TC(\textbf{X})=D_{KL}(P(X_{1},X_{2},...X_{N})||P(X_{1})\times P(X_{2})\times...\times P(X_{N}))$ (39)
@@ II.4.3 Co-Information
Where $\xi$ indicates all possible subsets of X. Unfortunately, unlike the total and dual total correlations, which easily accommodate large numbers of variables without changing the interpretation, the co-information can be hard to interpret for more than 3 elements. A significant issue is that the co-information can be negative, which has historically limited its applicability outside of theoretical contexts. [53] showed that, for three variables, the sign of co-information can be understood as indicating whether the statistical dependencies within the dataset are dominated by redundancy or synergy, although this pattern does not necessarily generalize (compare with the O-information, introduced in Sec. VI).
@@ III Information Dynamics
To begin with, what does it mean for a system to “compute”? This inevitably raises the complex, philosophical question “what is computation”, which we will sidestep entirely by operationalizing “computation” in a complex system as the process by which the system decides its immediate next state [64]. This can happen at every scale: individual elements can be in multiple states and decide their immediate next state based on the interactions between their own internal dynamics and the states of all their neighbors. At the macro-scale, the dynamics of the joint-state of the whole system “emerges” from the interactions of all it’s constituent elements [65].
@@
    Information storage [66], which characterizes how the past of an element or system constrains its immediate future,
@@ III.1 Conditional Entropy Rate
where $k$ indicates the number of discrete past states to take into account. Following the usual formalism, $X_{t}$ refers to the state of $X$ at time $t$, while $X_{-k:t-1}$ indicates a vector of states, stretching from time $t-1$ all the way back to time $t-1-k$. This tells us how surprising any particular realization $x$ is, given that it is preceded by the particular sequence $x_{-k:t-1}$. When $h_{\mu}(x)$ is high, then X has done something very unexpected and inconsistent with its past behavior, while when $h_{\mu}(x)$ is low, whatever state $X$ has just adopted would be unsurprising to anyone who had been watching it for the last $k$ timesteps. The local conditional entropy rate has been used in thermodynamic interpretations of the transfer entropy measure [75].
@@ III.2 Information Storage
The conditional entropy rate quantifies the “left over” uncertainty intrinsic to $X$ after the past has been accounted for. Its dual measure is the active information storage [66, 73], which quantifies how much information the past of a variable discloses about its immediate next state. In this context, “information storage” can be thought of as a measure of how dependent an element of a system is on it’s own past. If every next state is chosen at random, then $H_{\mu}(X)=H(X)$ and $A(X)=0$. Formally, active information storage is given by:
@@
Intuitively, this shows us that some of our uncertainty about the next state of $X$ could be resolved by looking at $X$’s past, and a residual part that cannot be resolved. This is a key insight when attempting to model complex systems. For processes with a naturally high entropy rate, simply observing the past behavior will not help predict the future (and consequently, modeling is hard). Conversely, a highly predictable periodic process will have high active information storage and the future can be easily predicted from the past. Importantly, both of these measures (and indeed, all of the basic information dynamics) are purely observational; they do not provide any insight into the “causal” structure of th system, and as such cannot be used to predict the effects of interventions or perturbations.
Active information storage has been applied in a wide range of contexts, from understanding distributed processing in neural networks [73], predictive coding models [76], eye and gaze tracking [77], to understanding the dynamics of football matches [78].
@@
This allows a “time-resolved” picture of how a single variable remembers and computes with its own past. When $a(x)>0$, the system is behaving “as we would expect”, in that past states are reducing our uncertainty about future states in accordance with the overall statistics of the system. In contrast, when $a(x)<0$, we are being “misinformed” and the system appears to be transiently “breaking” from it’s own past to do something unexpected.
@@ III.3 Information Transfer
Transfer entropy is probably the most widely used information dynamic in the study of complex systems, having been used in studies of biological networks of neurons, [84, 85, 86, 87, 88, 89], cryptocurrencies [90, 91, 92, 93], precious metal economics [94], the interaction between social media sentiment and commodity prices [93], the directed relationship between atmospheric carbon dioxide levels and global warming [95], climate models [96, 30, 97] and more (for further citations, see [83]). Transfer entropy’s nonparametric nature makes it particularly suited to analysis of complex systems where non-linearities are known to play a significant role in the dynamics.
@@ III.3.1 Conditional Transfer Entropy
The effect of conditioning on $Z$’s past gives us the amount of information $X$’s past gives us about $Y$’s future above and beyond that which is provided by $Y$’s own past and $Z$’s past. As with conditional mutual information, conditional transfer entropy is essential when assessing information dynamics in complex systems, which by definition have large numbers of interacting elements. For two variables $X$ and $Y$ belonging to a set of elements V, the complete transfer entropy between $X$ and $Y$ ($TE(X\to Y|\textbf{V}^{-XY})$) is given by conditioning $TE(X\to Y)$ on the past of every other element in V simultaneously (excluding $X$ and $Y$, of course). The conditional transfer entropy plays a key role in the discovery of effective connectivity networks from time-series data (see Sec. V.2 and [106, 107]). If conditioning $TE(X\rightarrow Y)$ on $\textbf{V}^{-XY}$ increases the overall transfer entropy, then $X$ and $Y$ share synergistic dependency that is only “illuminated” when all the other elements are accounted for. In contrast, if conditioning reduces the transfer entropy, then there are redundancies that are being conditioned out. For more on this, see [108]. Conditional transfer entropy can be localized in the same way as the standard bivariate transfer entropy.
@@ IV.1 PID Intuition
Where $Red(X_{1},X_{2};Y)$ corresponds to the information about $Y$ that is redundantly present in both $X_{1}$ and $X_{2}$ (i.e. could be learned by observing either $X_{1}$ alone or $X_{2}$ alone). $Unq(X_{1};Y/X_{2})$ is information about $Y$ that is uniquely present in $X_{1}$ (i.e. information that could only be learned by observing $X_{1}$), and likewise for $Unq(X_{2};Y/X_{1})$. Finally, $Syn(X_{1},X_{2};Y)$ is that information about $Y$ that is synergistically present in the joint states of $X_{1}$ and $X_{2}$ (and could not be learned by examining either $X_{1}$ or $X_{2}$ individually). A moments reflection should convince the reader that this enumerates all possible combinations.
@@
In general, however, for $N>2$, it is not the case that the relationship between mutual information terms and partial information “atoms” is so tightly constrained. For $N=3$, there are seven known mutual information terms but eighteen unique partial information atoms. To solve the general care involves developing some mathematical machinery beyond that which is provided by classical Shannon information theory.
@@ IV.2 Redundancy Lattices & Functions
Suppose we have a system of $N$ predictor variables, all informing about the state of a single target $Y$. We would like decompose the joint mutual information $I(X_{1},\ldots,K_{N};Y)$ down into all relevant combinations of sources. Doing this requites two things: the first is an operational definition of what it means for information to be “redundant” to two or more sources (which we will denote as $I_{\cap}$). The second is an understanding of the structure of multivariate information (which is given by the redundancy lattice described below). Formally, given some (as yet undefined) set of all meaningful sources of information ($\mathfrak{A}$), we want a decomposition such that:
@@
Where $I_{\partial}$ quantifies the information about $Y$ that could only be learned by observing $\boldsymbol{\alpha}$ and no simpler combination of sources. In the case of the two-input example given above, the sources $\boldsymbol{\alpha}$ are the redundant, two unique, and pure synergy terms.
@@ IV.2.1 The Redundancy Lattice
When decomposing $I(X_{1},\ldots,X_{N};Y)$, the first step is to enumerate the set of all possible sources of information. For example, information can be disclosed by any $X_{i}$, but it can also be disclosed by any pair {$X_{i},X_{j}\}$, every triple, and so on. The set S of all possible (potentially multivriate) sources is given by the power set of the set of $\{X_{i}\}$: $\textbf{S}=\mathcal{P}_{0}(\{X_{1},\ldots,X_{N}\})$, where the 0 subscript indicates that the empty set is being excluded from the power set (since a set of no elements discloses no information).
@@
The four elements of $\mathfrak{A}$ correspond to our intuitive ideas of redundant, unique, and synergistic information described above. The partial ordering shows that $\{X_{1}\}\{X_{2}\}\prec\{X_{1}\}$ and $\{X_{2}\}$, while both $\{X_{1}\}$ and $\{X_{2}\}\prec\{X_{1},X_{2}\}$. This defines the redundancy lattice shown in Fig. 4.
@@ IV.3 Information Modification, Revisited
This shows that the transfer entropy consists of two, non-overlapping informational components. The first is referred to as the state-independent information transfer (SITE): $Unq(X_{-k:t-1};Y_{t})$ is the information that passes from $X$’s past of $Y$’s future and is independent of the state of $Y$’s own past. In contrast, the state-dependent information transfer (SDTE), $Syn(X_{-k:t-1},Y_{-j:t-1};Y_{t})$, which is the novel information generated at time $t$ in $Y$ by the interaction of $X$’s past and $Y$’s past.
This complicates transfer entropy’s status as a measure of information transfer, as it appears to conflate the true transfer of information with synergistic information modification [135]. As it stands, the significance of this has largely remained under-explored. Daube, Gross, and Ince explored the contribution of synergy to sensory EEG data analysis, finding that the apparent transfer entropy was largely reflecting the synergistic component (with the unique “true flow” component being negative) [136]. More recently [64] argued that the transfer entropy could be understood as conflating information transfer and information modification (or computation), with the SITE reflecting information transfer and the SDTE reflecting a kind of computation that $Y$ performs based on its own past and the past of $X$. This is one area in particular with a lot of potential to support future work.
@@ IV.4.3 Integrated Information Decomposition
which can be computed for continuous variables under Gaussian assumptions, but does not readily localize to individual realizations. Unfortunately, while the $I^{MMI}_{\Phi\cap}$ function has practical appeal, it has counter-intuitive behaviors that have led some to question its appropriateness [150]. In particular, it appears to conflate first-order autocorrelation and higher-order temporal synergy in potentially fatal ways. Other double redundancy functions have been proposed, including one generalizing the $i^{sx}_{\cap}$ function [147], one based on the $i^{ccs}_{\cap}$ function [71], and one based on the dependency lattice approach of James et al., [120, 70].
@@ V Network Inference
One of the foundational methods in complex systems science is the construction of network models from data [154, 155]. It is hard to overstate how fundamental networks are to the modern study of complex systems, appearing in everything from brain science [156] to sociology [157, 158], from ecology [159, 160] to economics [161]. In applications, networks generally break down into two broad classes: there are physical networks (where the nodes and edge are unambiguous, physical objects in the real world), and statistical networks (where the links between nodes reflects some statistical dependency between them). Physical networks are generally reasonably easy to infer the structure of. For example, the airline network is conceptually simple: nodes may be airports (or cities), and a directed edge exists between two nodes if there is a regular flight from one to another [162]. Another example might be a structural connectivity network map of the brain: using diffusion tensor imaging (DTI) it is possible to get an estimate of the white matter tracts in the brain that physically connect disparate brain regions [163].
In contrast, statistical networks are more abstract: rather than referring directly to some physical structure in the real world (like a white-matter tract, or airline route), the edges in a statistical network refer to some degree of statistical dependency between elements. In an information-theoretic context, these dependencies almost always take the form of some kind of mutual information: nodes are usually said to be connected if knowing something about the state of one node reduces our uncertainty about the state of another node. There are other “kinds” of statistical networks not based on mutual information (for example, dynamic causal modeling builds statistical networks using Bayesian model comparison [164, 165]), however for our purposes, we will focus on information-theoretic approaches. Within the category of statistical networks, we will describe two general classes: the “functional” connectivity network, which has no notion of temporal order and produces undirected networks, and the “effective” connectivity network, which does take time into account (if time is a relevant feature) and produces directed networks. Finally, we will discuss how the problem of synergies in data can complicate the very idea of a bivariate network model of a complex system and briefly touch on higher-order models such as hypergraphs and simplicial complexes (although a rigorous mathematical treatment of these topics is beyond the scope of this tutorial).